# Ultrafast electrical control of dipolaritonic circuits with a few femto-joul per bit power consumption

### v1

@@ Abstract
The next generation of photonic circuits will require on-chip, ultrafast, compact, and energy-efficient components in a scalable platform to power technologies like quantum and neuromorphic computing. Here we present a significant advancement in ultrafast electrical control of highly nonlinear hybrid light-matter quasi-particles, called waveguide exciton-dipolaritons, with extremely low power consumption. Our device, that was shown to perform as an optical transistor with an optical switching power as low as $\sim$1 fJ/operation, achieves a GHz-rate electrical modulation with a record-low total energy consumption $\sim$3 fJ/bit and a compact active area of down to 25 $\mu$m2. This work establishes waveguide-dipolariton platforms for scalable, electrically reconfigurable, ultra-low power photonic circuits for both classical and quantum computing and communication.
dipolar exciton polaritons, photonic computing, optical modulators, optical transistor, neuromorphic computing, quantum photonics
Introduction: Integrated photonic circuits are most likely the best platform for the rapidly developed fields of optical information processing [1, 2, 3, 4, 5], including classical and quantum simulators and computers [6, 7], as well as neuromorphic calculators [8, 9].The most crucial requirements for complex optically-based reconfigurable circuitry for neuromorphic and quantum calculations [8, 10], is the ability for a scalable, ultrafast electrical control of individual nodes with an ultra-low power consumption per node per operation, together with a strong nonlinear optical response.
Exciton polaritons (EP) [11], resulting from the strong coupling of confined photons and excitons in semiconductor heterostructures, are excellent building blocks for photonic circuits. They can be directly addressed and optically manipulated due to their photonic part, and have significant optical nonlinearities due to the matter-like, excitonic part. Recent demonstrations of ultrafast all-optical control of EP in microcavities, including optical switching [12, 13, 14, 15, 16, 12, 17, 13, 16, 18, 19], EP condensate-based photonic lattice simulators [6, 7, 20], and EP based neuromprphic computing [9, 21] are just a few examples of the promise of EP platforms.
@@
For a truly complex photonic circuitry, electrically controlled EP-based nonlinear nodes should be densely integrated into monolithic chip geometries. For microcavity-based EP elements, only optical control has been demonstrated so far, which is inherently limited by scalability and optical loss [22]. Furthermore, significant optical power is required per node, typically more extensive than $10^{3}$/node/operation at 1GHz [12], and densely scaling up layers of nodes for deep circuits is a major challenge as microcavities cannot be easily stacked.
A more natural geometry for EP-based circuits is that of EP in chip-integrated waveguides, known as waveguide exciton-polaritons (WEP)[23, 24], for which various linear and non-linear photonic elements have been demonstrated recently in low loss devices [25, 26, 27, 28, 29, 30]. These recent demonstrations set WEP as a promising on-chip platform for complex photonic circuitry.
An important step towards locally reconfigurable and scalable WEP circuits is the recent realization of electrically controlled elements based on electrically gated dipolar WEP (DWEP). In such devices, wide quantum wells (QW) are embedded inside an optical waveguide. With top electrical gates at desired locations, tunable electric fields can be applied perpendicular to the QW plane and to the WEP propagation direction. These fields induce a quantum confined Stark effect on the WEP through their excitonic component, leading to an energy shift of the whole DWEP dispersion under the gate [24], allowing electrical control of the DWEP propagation. Furthermore, the local electric field polarizes the DWEP, resulting in WEP having an induced electric dipole. These DWEP then interact very strongly with each other, leading to a polariton record-high effective nonlinearity that is up to two-orders of magnitude enhanced as compared to unpolarized EP [28]. This huge enhancement is explained via a combination of direct dipole-dipole interactions and dipolar screening of the applied field [29], together with the reduction of correlations due to the ultralight mass of the DWEP [31]. The unique properties of such DWEP were utilized to demonstrate an electrically controlled photon switch, as is illustrated in Fig. 1, and a strongly nonlinear optical transistor, both operating with only few hundred photons, very low switching voltages of $\sim$1V, and small footprints of $\sim$100 $\mu$m2 [29]. Recently, electrically tunable quantum correlations and a partial 2-photon blockade was demonstrated, showing potential for realizing DWEP for electrically switchable, universal 2-photon gates [32]. If such photonic-building blocks can be reconfigured and operated at ultrafast rates, and can be shown to be scalable in terms of electrical power consumption and size, they will open up a route towards universal photonic computation on-chip.
In this work, we demonstrate a fast, sub-nanosecond temporal control of such nonlinear DWEP-switches operating at record-low powers of sub 3 Femto-Joule/bit and a very small footprint of $<50\mu m^{2}$ per node, a significant step towards complex optical circuitry based on WEP.
@@
Table I: Integrated electro-optic modulators Mechanism Material Power consumption ER Footprint [$fJ/bit$] @ 1 GHz  [^1] The power consumption was rescaled to a bit rate of 1GHz with duty-cycle of 0.5 [dB] [$\mu m^{2}$] DWEP (this work) GaAs 3similar-toabsent3$\sim$ 3∼ 3222see the SM for a full derivation 25333See DC measurements of ER in Fig. 3, limited by the SNR 25444projected Pockels effect (MZI [^5] Mach-Zender Interformeter) [33] LN 26 30 $10^{4}$ Carrier injection [34]  [^6] The device with the lowest power consumption Si 500 8 $10^{3}$ Plasmonic–organic–hybrid (MZI) [35] Si and Au 2750 25 777loss of 5dB per element $\sim 10$ MC EP (all optical) [12] GaAs [^8] Can also be realized in many other material systems, see Ref. [14] 1980 12 $1.6\times 10^{3}$
Electrical modulation of DWEP: to test the frequency response of the system, we injected WEP resonantly through the input grating using a CW laser, while modulating the gate voltage using a square electrical pulse $f(t)$ with a nanosecond rise and fall times. The experiment was repeated with different periods and duty cycles, as depicted schematically in Fig. 2(a). The field amplitudes and offsets of the electrical modulation were selected based on the DC measurements presented in Fig. 1 (f-h), and can be represented by $V(t)=V_{G}^{off}+\Delta V_{G}\times f(t)$. The output signal was imaged onto a streak camera. Such Streak images for two different modulation frequencies are presented in Fig. 2(b,c). Fig. 2(d,e) presents the time-domain DWEP transmission of single pulses (blue dots) plotted on top of the modulating electric field (solid green line) for WEPs excited at $-13.7,-4meV$, corresponding to excitonic fractions of $|\chi_{X}|^{2}=0.26,0.77$ respectively. The black lines indicate the “on” (solid) and “off” (dashed) states. The transmission data is fitted with a pulse function (red line):
@@
Bandwidth: From such fitting as above, we extracted $\tau_{r}$, $\tau_{f}$, as well as the extinction ratio (ER), $ER=T_{on}/T_{off}$, for various excitation energies, for electrical pulses with a 4MHz carrier frequency, and pulse durations of $15-20$ ns. The extracted rise and fall times are plotted in Fig. 3(a,b), and are both smaller for lower energy DWEPs, corresponding to more photon-like polaritons. Rise times as short as $0.5$ns and fall times as short as $1$ ns are measured. As seen from Fig. 2(d,e), the optical signal of photon-like DWEPs follows the electrical pulse. This means that the electrical pulse generator still limits the rise and fall times and not intrinsically by the WEP system. Therefore, we conclude that our devices can operate at $>$GHz bandwidth.
@@
Extinction ratio: the extinction ratio is shown for various DWEP energies in Fig. 3(c,d). Again, more photon-like DWEPs tend to have a higher extinction ratio, indicating more efficient switching. Importantly, these photon-dominated states also have faster propagation in the WG, lower propagation loss, and better in- and outcoupling efficiencies. $ER$ values as high as 15dB are measured, limited only by the SNR of the Streak camera. To demonstrate that the actual ER values are higher, we plot the $ER$ for the DC case, which had a better SNR of up to $25$ dB.
@@
Energy consumption: the energy consumption for our device is composed of two factors, the energy of photons in a cycle and the electrical energy to operate the switching, given by the current-voltage product in a cycle. To estimate this, we measured the current of the device during operation, with $V_{G}^{off}=3.1V$ and $\Delta V_{G}$ of 2.25, we measured currents of $I=$ 0.46 and 0.85 $\mu$A, respectively. For a duty cycle of $R=50\%$ and frequency of $f=1$ GHz, the energy consumption per operation is $1.3fJ/bit$; the measurement details appear in the SM. Optically, the optical power required for the nonlinear DWEP-transistor operation is 1.2 fJ per cycle at 1 GHz is required; (see SM). Finally, electrical and optical energy consumption adds up to about $2.5fJ$ per cycle at 1 GHz for a fully functional electrically-modulated transistor device.
@@
Discussion: We demonstrated DWEP signal modulation with a bandwidth exceeding 1 GHz, primarily constrained by the limitations of the electrical pulse generator and the electronic circuitry, together with a very high extinction ratio. A proper electrical design can significantly improve the bandwidth, as was demonstrated in GaAs-based electro-optical devices that have achieved modulation frequencies up to $100GHz$ [36]. With a minimal footprint (down to 25 $\mu m^{2}$), such elements are well suited for high-density photonic circuitry. The footprint could be further reduced by shortening the gate length to a few times the WEP wavelength in the medium ($\lambda\simeq 240nm$), and decreasing the channel width to $\sim 0.5\mu m$ by side etching [25, 27], resulting in an overall footprint as small as $\sim 1\mu m^{2}$ per node.
@@
Finally, the same device displays an electrically tunable, optical nonlinearity which is, as far as we know, the highest of any exciton-polariton-based system, and is comparable to that of Rydberg atoms [32]. As mentioned before, this led to a demonstration of a fully operational optical transistor operating with 1.2 fJ/bit [29], and to quantum correlations at the two-photon level [32], which can be controlled very accurately by the applied electric field. The current demonstration of local >GHz modulation of an optical node, will allow a reconfigurable polariton-based optical circuitry, where the linear and nonlinear function at each node can be controlled separately and be reconfigured each operation, making electrically polarized DWEP as excellent candidates for reconfigurable deep optical circuits for either neuromorphic or quantum processors in a monolithic platform. Low propagation loss is also essential for a deep photonic circuit, particularly at the quantum limit where signal amplification is impossible. Our current device has a propagation loss of approximately 4-10 dB/cm[25], with a lower loss for high photonic fractions, which is a typical value for photonic losses in AlGaAs WG systems. The photonic loss can be significantly reduced to 0.4 dB/cm by integrating AlGaAs on an insulator (AlGaAsOI) [38]. Given the achievable element density of AlGaAsOI, the effective loss per number of elements compares with state-of-the-art silicon-on-insulator (SOI) systems. Finally, for full integration, the chip should also incorporate single-photon detectors, which have already been realized in similar AlGaAs systems using superconductor-nanowire single-photon detectors (SNSPDs) [5, 39].
Acknowledgements
R.R. and D.L. acknowledge the support from the Israeli Science Foundation Grant 1087/22, and from the NSF-BSF Grant 2019737. H.D. acknowledges the support of the National Science Foundation under grant DMR 2004287, the Army Research Office under grant W911NF-17-1-0312, and the Air Force Office of Scientific Research under grant FA2386-21-1-4066, and the Gordon and Betty Moore Foundation under grant N031710. This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF9615 to L. N. Pfeiffer, and by the National Science Foundation MRSEC grant DMR 2011750 to Princeton University

### v2

# Ultrafast electrical control of dipolariton-based optical circuits with a few femto-joul per bit power consumption

@@ Abstract
The next generation of photonic circuits will require programmable, ultrafast, and energy-efficient components on a scalable platform for quantum and neuromorphic computing. Here, we present ultrafast electrical control of highly nonlinear light-matter hybrid quasi-particles, called waveguide exciton-dipolaritons, with extremely low power consumption. Our device performs as an optical transistor with a GHz-rate electrical modulation at a record-low total energy consumption $\sim$3 fJ/bit and a compact active area of down to 25 $\mu$m2. This work establishes waveguide-dipolariton platforms for scalable, electrically reconfigurable, ultra-low power photonic circuits for both classical and quantum computing and communication.
I Introduction
Integrated photonic circuits are the most promising platform for the rapidly developing fields of optical information processing [1, 2, 3, 4, 5], including classical and quantum simulators and computers [6, 7], as well as neuromorphic calculators [8, 9]. A crucial requirement of an integrated photonic platform [8, 10] is the ability of ultrafast electrical control of individual nodes with ultra-low power consumption per node per operation.
Exciton polaritons (EP) [11], resulting from the strong coupling of confined photons and two-dimensional excitons, are excellent building blocks for photonic circuits. They can be directly addressed due to their photonic part, and they have stronger optical nonlinearities due to their excitonic part. Recent demonstrations of ultrafast all-optical control of EP in microcavities, including optical switching [12, 13, 14, 15, 16, 12, 17, 13, 16, 18, 19], EP condensate-based photonic lattice simulators [6, 7, 20], and EP based neuromprphic computing [9, 21] are just a few examples of the promise of EP platforms.
@@ I Introduction
Yet, for modern complex photonic circuitry, highly desirable are electrically controlled photonic nodes that are densely integrated into monolithic chip geometries. For microcavity-based EP elements, only optical control has been demonstrated so far; significant optical power is required per node, typically exceeding $10^{3}$/node/operation at 1GHz [12]; scaling up layers of nodes for deep circuits is also a major challenge, as microcavities cannot be easily stacked.
Alternatively, monolithic chip-integrated waveguide exciton-polaritons (WEP) [22, 23] are well suited for scalable EP circuits. Several key elements have been demonstrated recently [24, 25, 26, 27, 28, 29]. An important step toward locally reconfigurable and scalable waveguide exciton-polariton (WEP) circuits is the realization of electrically gated dipolar WEP (DWEP) structures, where a top gate applies a perpendicular electric field to wide quantum wells (QWs) inside the waveguide, inducing a quantum-confined Stark effect and voltage-controlled exciton dipoles [23]. These dipoles exhibit strong dipole–dipole interactions and screening effects, enhanced further by the ultralight effective mass of the polaritons [30], resulting in record-high effective nonlinearities—up to two orders of magnitude larger than in unpolarized polaritons [27]—and enabling key demonstrations such as electrically controlled few-photon transistors [28]. Remarkably, electrically tunable quantum correlations and a partial 2-photon blockade were recently demonstrated, showing potential for realizing DWEP for electrically switchable, universal 2-photon gates [31].
Here, we demonstrate a fast, sub-nanosecond temporal control of such nonlinear DWEP-switches operating at record-low powers of sub 3 Femto-Joule/bit and a very small footprint of $<50\mu m^{2}$ per node, a significant step towards complex optical circuitry based on WEP, towards universal photonic computation.
@@ II DWEP device characterization
Table I: Integrated electro-optic modulators Mechanism Material Power consumption ER Footprint [$fJ/bit$] @ 1 GHz [dB] [$\mu m^{2}$] DWEP (this work) GaAs $\sim 3$∗∗ 25∗∗∗ 25 Pockels effect (MZI††) [32] LN 26 30 $10^{4}$ Carrier injection [33] Si 500 8 $10^{3}$ Plasmonic–organic–hybrid (MZI) [34] Si and Au 2750 25 111loss of 5dB per element $\sim 10$ MC EP (all optical) [12] GaAs‡‡ 1980 12 $1.6\times 10^{3}$
The power consumption was rescaled to a bit rate of 1GHz with duty-cycle of 0.5
∗∗ see the SM for a full derivation
∗∗∗ See DC measurements of ER in Fig. 3, limited by the SNR
projected
†† Mach-Zender Interformeter
The device with the lowest power consumption
‡‡ Can also be realized in many other material systems, see Ref. [14]
III Electrical modulation of DWEP
to test the frequency response of the system, we injected WEP resonantly through the input grating using a CW laser, while modulating the gate voltage using a square electrical pulse $f(t)$ with a nanosecond rise and fall times. The experiment was repeated with different periods and duty cycles, as depicted schematically in Fig. 2(a). The field amplitudes and offsets of the electrical modulation were selected based on the DC measurements presented in Fig. 1 (f-h), and can be represented by $V(t)=V_{G}^{off}+\Delta V_{G}\times f(t)$. The output signal was imaged onto a streak camera. Such Streak images for two different modulation frequencies are presented in Fig. 2(b,c). Fig. 2(d,e) presents the time-domain DWEP transmission of single pulses (blue dots) plotted on top of the modulating electric field (solid green line) for WEPs excited at $-13.7,-4meV$, corresponding to excitonic fractions of $|\chi_{X}|^{2}=0.26,0.77$ respectively. The black lines indicate the “on” (solid) and “off” (dashed) states. The transmission data is fitted with a pulse function (red line):
@@ III Electrical modulation of DWEP
From such fitting as above, we extracted $\tau_{r}$, $\tau_{f}$, as well as the extinction ratio (ER), $ER=T_{on}/T_{off}$, for various excitation energies, for electrical pulses with a 4MHz carrier frequency, and pulse durations of $15-20$ ns. The extracted rise and fall times are plotted in Fig. 3(a,b), and are both smaller for lower energy DWEPs, corresponding to more photon-like polaritons. Rise times as short as $0.5$ns and fall times as short as $1$ ns are measured. As seen from Fig. 2(d,e), the optical signal of photon-like DWEPs follows the electrical pulse. This means that the electrical pulse generator still limits the rise and fall times and not intrinsically by the WEP system. Therefore, we conclude that our devices can operate at $>$GHz bandwidth.
Extinction ratio:
@@
the extinction ratio is shown for various DWEP energies in Fig. 3(c,d). Again, more photon-like DWEPs tend to have a higher extinction ratio, indicating more efficient switching. Importantly, these photon-dominated states also have faster propagation in the WG, lower propagation loss, and better in- and outcoupling efficiencies. $ER$ values as high as 15dB are measured, limited only by the SNR of the Streak camera. To demonstrate that the actual ER values are higher, we plot the $ER$ for the DC case, which had a better SNR of up to $25$ dB.
Energy consumption:
@@
the energy consumption for our device is composed of two factors, the energy of photons in a cycle and the electrical energy to operate the switching, given by the current-voltage product in a cycle. To estimate this, we measured the current of the device during operation, with $V_{G}^{off}=3.1V$ and $\Delta V_{G}$ of 2.25, we measured currents of $I=$ 0.46 and 0.85 $\mu$A, respectively. For a duty cycle of $R=50\%$ and frequency of $f=1$ GHz, the energy consumption per operation is $1.3fJ/bit$; the measurement details appear in the SM. Optically, the optical power required for the nonlinear DWEP-transistor operation is 1.2 fJ per cycle at 1 GHz is required; (see SM). Finally, electrical and optical energy consumption adds up to about $2.5fJ$ per cycle at 1 GHz for a fully functional electrically-modulated transistor device.
IV Discussion
@@
We demonstrated DWEP signal modulation with a bandwidth exceeding 1 GHz, primarily constrained by the limitations of the electrical pulse generator and the electronic circuitry, together with a very high extinction ratio. A proper electrical design can significantly improve the bandwidth, as was demonstrated in GaAs-based electro-optical devices that have achieved modulation frequencies up to $100GHz$ [35]. With a minimal footprint (down to 25 $\mu m^{2}$), such elements are well suited for high-density photonic circuitry. The footprint could be further reduced by shortening the gate length to a few times the WEP wavelength in the medium ($\lambda\simeq 240nm$), and decreasing the channel width to $\sim 0.5\mu m$ by side etching [24, 26], resulting in an overall footprint as small as $\sim 1\mu m^{2}$ per node.
@@
Finally, the same device displays an electrically tunable, optical nonlinearity which is, as far as we know, the highest of any exciton-polariton-based system, and is comparable to that of Rydberg atoms [31]. As was mentioned before, this led to a demonstration of a fully operational optical transistor operating with 1.2 fJ/bit [28], and to quantum correlations at the two-photon level [31], which can be controlled very accurately by the applied electric field. The current demonstration of local >GHz modulation of an optical node, will allow a reconfigurable polariton-based optical circuitry, where the linear and nonlinear function at each node can be controlled separately and be reconfigured for each operation, making electrically polarized DWEP an excellent candidates for reconfigurable deep optical circuits for either neuromorphic or quantum processors in a monolithic platform. Low propagation loss is also essential for a deep photonic circuit, particularly at the quantum limit where signal amplification is impossible. Our current device has a propagation loss of approximately 4-10 dB/cm[24], with a lower loss for high photonic fractions, which is a typical value for photonic losses in AlGaAs WG systems. The photonic loss can be significantly reduced to 0.4 dB/cm by integrating AlGaAs on an insulator (AlGaAsOI) [37]. Given the achievable element density of AlGaAsOI, the effective loss per number of elements compares with state-of-the-art silicon-on-insulator (SOI) systems. Finally, for full integration, the chip should also incorporate single-photon detectors, which have already been realized in similar AlGaAs systems using superconductor-nanowire single-photon detectors (SNSPDs) [5, 38].
Funding.
R.R. and D.L. acknowledge the support from the Israeli Science Foundation Grant 1087/22, and from the NSF-BSF Grant 2019737. H.D. acknowledges the support of the National Science Foundation under grant DMR 2004287, the Army Research Office under grant W911NF2510055, the Air Force Office of Scientific Research under grant FA2386-21-1-4066, and the Gordon and Betty Moore Foundation under grant N031710. This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF9615 to L. N. Pfeiffer, and by the National Science Foundation MRSEC grant DMR 2011750 to Princeton University